What's the answer for :
3x + y = 1
x - 2y = 5

Answer by unlockmath(1688)

Hello,
Let's start by multiplying the first equation by 2:
6x + 2y = 2
x - 2y = 5
Now we can add these to get:
7x=7 Divide both sides by 7 will be:
x=1
Plug in 1 into the first equation and we get:
6+2y=2 Subtract 6 from both sides:
2y=-4 Divide by 2:
y=-2
There we go the answers are x=1 and y=-2
Plug these into both equations and see what happens.
Make sense?
RJ
Check out a new book at:
www.math-unlock.com